OpenStudy (anonymous):

Simplify each sum or difference 4sqrt{7}-sqrt{63}

OpenStudy (radar):

\[4\sqrt{7}-\sqrt{63}=4\sqrt{7}-\sqrt{9}\sqrt{7}=4\sqrt{7}-3\sqrt{7}=\sqrt{7}\]

OpenStudy (anonymous):

So the simplified answer would be \[\sqrt{7}\] ?